Part A How much energy can be obtained from conversion of 1.8 gram of mass? Express...

Part A

How much energy can be obtained from conversion of 1.8 gram of mass?

Express your answer to two significant figures and include the appropriate units.

E =

SubmitMy AnswersGive Up

Part B

How much mass could this energy raise to a height of 1.5 km above the Earth's surface?

Express your answer to two significant figures and include the appropriate units.

m =
Answer #1

A)

given

m=1.8g =0.0018kg

velocity of light c= 3*10^8

energy = mc^2

E =0.0018*300000000*300000000

E =1.62*10^{14}J

E =162000Gj

B)

given

h=1.5km =1500m

E=m*h*g

m=E/(h*g)

m=162000/1500*9.8

m=162000/14700

m=11.02 kg